Please explain the below problem and show the calcuation correct answers are at the end of the problem.

The Larkspur Furniture Company's chair manufacturing requires multiple shades of wood stain. The most popular stain color is "Warm Oak”. The company's production operation uses 560 cases per year of this stain. The ordering cost is $43 per order. The company estimates the holding costs at $8 per case per year. The lead time to receive a shipment from the supplier is 5 days. The company operates 250 days per year.  

What is the optimal number of cases to order at a time?

This problem does not require you to enter the units of measure for your answers. Don't type anything in the response box below except the final calculation result.  

Answer:

The correct answer is: 78

Question text

The Larkspur Furniture Company's chair manufacturing requires multiple shades of wood stain. The most popular stain color is "Warm Oak”. The company's production operation uses 560 cases per year of this stain. The ordering cost is $43 per order. The company estimates the holding costs at $8 per case per year. The lead time to receive a shipment from the supplier is 5 days. The company operates 250 days per year.  

What is the average number of cases on hand?

This problem does not require you to enter the units of measure for your answers. Don't type anything in the response box below except the final calculation result.  

Answer:

Average Inventory Level = Q*/2

The correct answer is: 39.00

Question text

The Larkspur Furniture Company's chair manufacturing requires multiple shades of wood stain. The most popular stain color is "Warm Oak”. The company's production operation uses 560 cases per year of this stain. The ordering cost is $43 per order. The company estimates the holding costs at $8 per case per year. The lead time to receive a shipment from the supplier is 5 days. The company operates 250 days per year.  

How many orders per year will there be for this inventory item?

This problem does not require you to enter the units of measure for your answers. Don't type anything in the response box below except the final calculation result.  

Answer:

N = D/Q*

The correct answer is: 7.18
Answer #1

Question 1:

To be able to do this question, you must understand the concept of economic order quantity (EOQ). EOQ is number of units to be ordered per order, so that the company incurs minimum cost of ordering and storing the units. At EOQ the total ordering cost and total storage (carrying) cost is equal. This is the most important point regarding EOQ. The formula for EOQ is as given below:

EOQ= square root of [(2*A*O)/C]

here,

A= the annual demand for the year, given to be 560 cases

O= ordering cost per order, given as $43 per order

C= carrying cost per unit per year, given as $8 per order

EOQ= square root of [(2*560*43)/8]

=Square root of [48160/8]

=Square root of [6020]

=78 units (rounded off)

Question 2:

Thus question relates to average inventory. The formula for average inventory is as follows:

Average Inventory= safety stock+ [(re order quantity)/2]

Reorder quantity means the number of units ordered per order. We have chosen to place orders at EOQ. Therefore the reorder quantity corresponds to EOQ in this case. Also there is no mention of safety stock in this question, therefore safety stock=0. The above formula transforms to:

Average Inventory= 0+ [(EOQ/2)]

= EOQ/2 = 78/2= 39 units

Question 3:

To compute the number of orders we need to have information regarding the number or units required in the year. The yearly requirement is given to be 560 cases. We have chosen to order at EOQ ie, 78 units per order. Now, to find how many orders it takes to meet the annual requirement, we have to divide the annual requirement by the number of units per order as shown below:

Number of orders= annual requirement/ reorder quantity

= 560/78= 7.18(rounded off to two decimals)